### v1

# Computing the $q$-Multiplicity of the Positive Roots of $\mathfrak{sl}_{r+1}(\mathbb{C})$ and Products of Fibonacci Numbers

Abstract
Using Kostant’s weight multiplicity formula, we describe and enumerate the terms contributing a nonzero value to the multiplicity of a positive root $\mu$ in the adjoint representation of $\mathfrak{sl}_{r+1}(\mathbb{C})$, which we denote $L(\tilde{\alpha})$, where $\tilde{\alpha}$ is the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$. We prove that the number of terms contributing a nonzero value in the multiplicity of the positive root $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}$ with $1\leq i\leq j\leq r$ in $L(\tilde{\alpha})$ is given by the product $F_{i}\cdot F_{r-j+1}$, where $F_{n}$ is the $n^{\text{th}}$ Fibonacci number. Using this result, we show that the $q$-multiplicity of the positive root $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}$ with $1\leq i\leq j\leq r$ in the representation $L(\tilde{\alpha})$ is precisely $q^{r-h(\mu)}$, where $h(\mu)=j-i+1$ is the height of the positive root $\mu$. Setting $q=1$ recovers the known result that the multiplicity of a positive root in the adjoint representation of $\mathfrak{sl}_{r+1}(\mathbb{C})$ is one.
1 Introduction
Let $G$ be a simple linear algebraic group over $\mathbb{C}$, $T$ be a maximal algebraic torus in $G$ of dimension $r$, and $B$ be a choice of a Borel subgroup where $T\subseteq B\subseteq G$. Let $\mathfrak{g}$, $\mathfrak{h}$, and $\mathfrak{b}$ be the Lie algebras of $G$, $T$, and $B$, respectively. Let the set of roots corresponding to $(\mathfrak{g},\mathfrak{h})$ be denoted as $\Phi$ and let the set of positive roots with respect to $\mathfrak{b}$ be denoted as $\Phi^{+}$. Note that $\Phi^{+}\subseteq\Phi$. Let the simple roots be denoted as $\Delta\subseteq\Phi^{+}$. We let the Weyl group corresponding to $G$ and $T$ be denoted as $W=Norm_{G}(T)/T$. Recall that the Weyl group is generated by reflections about hyperplanes lying perpendicular to the simple roots of the Lie algebra $\mathfrak{g}$. For any $\sigma\in W$, the length of $\sigma$, denoted $\ell(\sigma)$, is equal to the minimal nonnegative integer $k$ such that $\sigma$ is a product of $k$ simple reflections.
It is well-known that a finite-dimensional irreducible representation $V$ of a simple Lie algebra $\mathfrak{g}$ can be studied by decomposing
$V=\oplus V_{\alpha}$ (1)
where the direct sum is indexed by a finite set of weights, which are linear functionals from the Cartan subalgebra to $\mathbb{C}$. When given a weight $\mu$, the corresponding subspace $V_{\mu}$ is called a weight space and the dimension of $V_{\mu}$ is called the multiplicity of $\mu$. To study representations of $\mathfrak{g}$, it is sufficient to determine the multiplicity of the weights appearing in equation (1).
To compute the multiplicity of a weight $\mu$ in the irreducible representation with highest weight $\lambda$, we can use Kostant’s weight multiplicity formula [10]:
$m(\lambda,\mu)=\sum_{\sigma\in W}(-1)^{\ell(\sigma)}\wp(\sigma(\lambda+\rho)-\mu-\rho)$ (2)
where $\rho=\frac{1}{2}\sum_{\alpha\in\Phi^{+}}\alpha$ and $\wp(\xi)$ is Kostant’s partition function. Recall that Kostant’s partition function $\wp(\xi)$ counts the number of ways the weight $\xi$ can be written as a nonnegative integral sum of positive roots.
When using equation (2), every element in the Weyl group contributes a term to the multiplicity. This forms a challenge since the order of the Weyl group increases factorially as the rank of the Lie algebra increases and many of the elements contribute trivially (a zero term) to the computation. In order to formulate a more efficient way to know what elements of the Weyl group contribute to Kostant’s weight multiplicity formula, we recall the following definition [6, Definition 1.1].
Definition 1.1.
For $\lambda,\mu$ integral weights of $\mathfrak{g}$, define the Weyl alternation set to be
$\mathcal{A}(\lambda,\mu)=\{\sigma\in W~{}|~{}\wp(\sigma(\lambda+\rho)-\mu-\rho)>0\}.$ (3)
The Weyl alternation set tells us the elements of the Weyl group for which $\sigma(\lambda+\rho)-\mu-\rho$ can be written as a nonnegative integral sum of positive roots.
The importance of the Weyl alternation set is that it reduces the multiplicity computation to its support. On this smaller contributing set, it is often the case that we can give formulas for the $q$-analog of Kostant’s partition function. We recall the definition next.
Definition 1.2.
The $q$-analog of Kostant’s partition function is the polynomial-valued function defined by
$\wp_{q}(\xi)=c_{0}+c_{1}q+c_{2}q^{2}+\cdots+c_{k}q^{k},$
where $c_{i}$ equals the number of ways to write $\xi$ as a sum of exactly $i$ positive roots.
The work on enumerating and characterizing the Weyl alternation sets for certain pairs of weights, $(\lambda,\mu)$, has gathered some attention in the literature. This includes the following. If $\lambda$ and $\mu$ are integral weights of the Lie algebras $\mathfrak{sl}_{3}(\mathbb{C})$ and $\mathfrak{sl}_{4}(\mathbb{C})$, then the sets $\mathcal{A}(\lambda,\mu)$ are known in general [3, 11]. Harris considered $\tilde{\alpha}$, the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$ with $r\geq 1$, and $\mu=0$ and established the cardinality of the Weyl alternation set, $|\mathcal{A}(\tilde{\alpha},0)|$, is a Fibonacci number [6, Theorem 1.2]. Harris also stated that if $\mu$ is a dominant weight, then the Weyl alternation set is either $\{1\}$ when $\mu=\tilde{\alpha}$ or empty otherwise. Chang, Harris, and Insko extended these results and showed that when $\mu=0$ and $\lambda$ is the sum of all the simple roots of the simple Lie algebra $\mathfrak{so}_{2r+1}(\mathbb{C})$, the cardinality of the Weyl alternation set $\mathcal{A}(\lambda,0)$ is given by a Fibonacci number [2, Corollary 2.1 and 3.1]. While in the Lie algebras $\mathfrak{sp}_{2r}(\mathbb{C})$ and $\mathfrak{so}_{2r}(\mathbb{C})$, the cardinality of the Weyl alternation sets $\mathcal{A}(\lambda,0)$ are given by a multiple of a Lucas number [2, Corollary 4.1 and 5.1]. Harris, Insko, and Williams consider $\tilde{\alpha}$, the highest root of the classical Lie algebras, and $\mu=0$ and show that $|\mathcal{A}(\tilde{\alpha},0)|$ satisfies homogeneous linear recurrence relations [8]. In the type $B$ case, if $\mu=\varpi_{1}$ denotes the first fundamental root, then $|\mathcal{A}(\lambda,\varpi_{1})|$ is a Fibonacci number [8, Corollary 8.4]. Harris et al. also consider the case where $\mu$ is a dominant positive root and give the multiplicity of these weights in the adjoint representation of the classical Lie algebras. Berenshteĭn and Zelevinskiĭ [1, Theorem 1.3] provided a comprehensive list of pairs of weights $(\lambda,\mu)$ of the classical Lie algebras and the exceptional Lie algebra $G_{2}$ for which $m(\lambda,\mu)=1$. Harris, Rahmoeller, Schneider, and Simpson consider certain such pairs, determine $\mathcal{A}(\lambda,\mu)$, and show that the $q$-multiplicity is a power of $q$ [9]. For all pairs of weights that Berenshteĭn and Zelevinskiĭ provided, Harris et. al conjecture that $m_{q}(\lambda,\mu)$ is a power of $q$ [9, Conjecture 7].
As noted above, past work fixed $\lambda$ to be the highest root or the sum of the simple roots of a classical Lie algebra, while $\mu$ was either 00 or a dominant positive root. In this work, we focus our study on $\mathfrak{sl}_{r+1}(\mathbb{C})$ and let $\Delta=\{\alpha_{1},\alpha_{2},\ldots,\alpha_{r}\}$ be the set of simple roots and $\Phi^{+}=\Delta\cup\{\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}:1\leq i<j\leq r\}$ be the set of positive roots. Throughout all of our results, we fix $r\geq 1$, let $\tilde{\alpha}$ be the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$, and consider $\mu\in\Phi^{+}$. In this way, $\mu$ may be a nondominant weight. Recall, that it is known that whenever $\mu$ is a root, $m(\tilde{\alpha},\mu)=1$. Hence, when $\mu$ is dominant, the pair $(\lambda,\mu)$ is given by Berenshteĭn and Zelevinskiĭ [1]. Moreover, we recall that the multiplicity is invariant under the action of the Weyl group. Namely, $m(\tilde{\alpha},\mu)=m(\tilde{\alpha},\sigma(\mu))$ for all $\sigma\in W$ [4, Proposition 3.1.20]. However, when $\mu\in\Phi^{+}$ is nondominant, a characterization and enumeration of the elements in $\mathcal{A}(\tilde{\alpha},\mu)$ as well as the value of $m_{q}(\tilde{\alpha},\mu)$ were unknown. Our work contributes this to the literature by proving the following results.
Theorem 2.1.
Let $r\geq 1$ and let $\tilde{\alpha}$ be the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$. Fix $1\leq i\leq j\leq r$ and let $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}\in\Phi^{+}$. Then $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$ if and only if $\sigma=1$ or $\sigma=s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}$ for some collection of nonconsecutive integers $12\leq\ell_{1},\ell_{2},\ldots,\ell_{k}\leq i-$ 1 and $j+1\leq\ell_{k+1},\ldots,\ell_{t}\leq r-1$.
From the characterization in Theorem 2.1, we are able to provide an enumeration for the cardinality of the Weyl alternation set.
Corollary 2.5.
Let $r\geq 1$ and let $\tilde{\alpha}$ be the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$. Fix $1\leq i\leq j\leq r$ and let $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}\in\Phi^{+}$. Then
$|\mathcal{A}(\tilde{\alpha},\mu)|=F_{i}\cdot F_{r-j+1},$
[…]
Given the characterization and enumeration of the terms contributing to the $q$-multiplicity formula, we are now able to provide the following result.
Theorem 4.1.
Let $r\geq 1$ and let $\tilde{\alpha}$ be the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$. Fix $1\leq i\leq j\leq r$ and let $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}\in\Phi^{+}$. Then
$m_{q}(\tilde{\alpha},\mu)=q^{r-h(\mu)},$
where $h(\mu)=j-i+1$ denotes the height of $\mu$.
Observe that setting $q=1$ in Theorem 4.1 recovers the known result that the multiplicity of a positive root in the adjoint representation is equal to one [4, Theorem 2.4.1].
This work is organized as follows: In Section 2, we present the proof to Theorem 2.1 showing that the cardinality of the Weyl alternation is a product of Fibonacci numbers. In Section 3, we establish results related to the $q$-multiplicity. In Section 4, we present the proof to Theorem 4.1 showing that the $q$-multiplicity of a positive root simplifies down to a power of $q$.
Acknowledgements. The author thanks her adviser Pamela E. Harris for introduction to the problem and feedback on an earlier draft. She also thanks Jeb Willenbring for insightful comments throughout the progression of this manuscript. The author gratefully acknowledge funding from the UW-Milwaukee Department of Mathematical Sciences in support of this research.
2 Weyl Alternation Sets
In [2], they consider the case where $\lambda$ is the highest root and $\mu=0$. We will still let $\lambda$ be the highest root and denote it as $\tilde{\alpha}$. However, we let $\mu\in\Phi^{+}$. The following result characterizes the elements in the set $\mathcal{A}(\tilde{\alpha},\mu)$.
Theorem 2.1.
$\sigma\in\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$ if and only if $\sigma=1$ or $\sigma=s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}$ for some collection of nonconsecutive integers $12\leq\ell_{1},\ell_{2},\ldots,\ell_{k}\leq i-$ 1 and $j+1\leq\ell_{k+1},\ldots,\ell_{t}\leq r-1$.
Proof.
($\Rightarrow$) We proceed by contradiction and assume $\sigma\neq 1$ or $\sigma\neq s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}$ for some collection of nonconsecutive integers $12\leq\ell_{1},\ell_{2},\ldots,\ell_{k}\leq i-$ 1 and $j+1\leq\ell_{k+1},\ldots,\ell_{t}\leq r-1$. Then $\sigma$ must be either $s_{1}$, $s_{r}$, $s_{y}$ with $i\leq y\leq j$, or must contain some product of $s_{n}$ and $s_{n+1}$ for some integer $1\leq n\leq r$.
Let $\sigma=s_{1}$. Then
$\wp(s_{1}(\tilde{\alpha}+\rho)-\rho-(\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}))$ $=\wp(\tilde{\alpha}-\alpha_{1}+\rho-\alpha_{1}-\rho-(\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}))$
$=\wp(\tilde{\alpha}-2\alpha_{1}-(\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}))$
$=\wp(-\alpha_{1}+\alpha_{2}+\cdots+\alpha_{i-1}+\alpha_{j+1}+\cdots+\alpha_{r})$
$=0$
since $-\alpha_{1}+\alpha_{2}+\cdots+\alpha_{i-1}+\alpha_{j+1}+\cdots+\alpha_{r}$ cannot be written as a sum of positive roots since the coefficient for $\alpha_{1}$ is negative. Thus, we get a contradiction, and $s_{1}\notin\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$. By [8, Proposition 3.4], if $\sigma\notin\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$, then neither is any $\sigma^{\prime}\in W$ that contains $\sigma$ in its reduced word expression. Thus, $\sigma\in W$ containing $s_{1}$ in its reduced word expression cannot be in $\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$. This reaches a contradiction to the assumption that $\sigma\in\mathcal{A}(\tilde{\alpha},\alpha_{i}+\cdots+\alpha_{j})$.
A similar argument is made for when $\sigma$ is either $s_{r}$, $s_{y}$ with $i\leq y\leq j$, or must contain some product of $s_{n}$ and $s_{n+1}$ for some integer $1\leq n\leq r$.
We have shown that if $\sigma\in\mathcal{A}(\tilde{\alpha},\alpha_{i}+\cdots+\alpha_{j})$, then $\sigma$ cannot contain $s_{1}$, $s_{r}$, $s_{k}$ with $i\leq k\leq j$, nor any product of consecutive $s_{n}$ and $s_{n+1}$ with $11\leq n\leq r-$ 1. Thus, the only remaining possibilities are that $\sigma=1$ or that $\sigma=s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}$ for some collection of nonconsecutive integers $12\leq\ell_{1},\ell_{2},\ldots,\ell_{k}\leq i-$ 1 and $j+1\leq\ell_{k+1},\ldots,\ell_{t}\leq r-1$ as desired.
($\Leftarrow$) It suffices to consider the following two cases.
Case 1: If $\sigma=1$, then
$\wp(1(\tilde{\alpha}+\rho)-\rho-(\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}))$ $=\wp(\tilde{\alpha}-(\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}))$
$=\wp(\alpha_{1}+\cdots+\alpha_{i-1}+\alpha_{j+1}+\cdots+\alpha_{r})$
$>0$
since $\alpha_{1}+\cdots+\alpha_{i-1}+\alpha_{j+1}+\cdots+\alpha_{r}$ can be written as a sum of positive roots. Thus,
$\sigma=1\in\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$.
Case 2: Let $\sigma=s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}$ where for some collection of nonconsecutive integers
$12\leq\ell_{1},\ell_{2},\ldots,\ell_{k}\leq i-$ 1 and $j+1\leq\ell_{k+1},\ldots,\ell_{t}\leq r-1$. Then
$\wp(s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}$ $(\tilde{\alpha}+\rho)-\rho-(\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}))$
$=\wp\left(\tilde{\alpha}+\rho-\sum_{h=1}^{t}\alpha_{\ell_{h}}-\rho-(\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})\right)$
$=\wp\left(\tilde{\alpha}-(\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})-\sum_{h=1}^{t}\alpha_{\ell_{h}}\right).$ (4)
Notice that $\{i,i+1,\ldots,j\}\cap\{\ell_{1},\ell_{2},\ldots,\ell_{t}\}=\emptyset$. Now if $x\in\{i,i+1,\ldots,j\}\cup\{\ell_{1},\ell_{2},\ldots,\ell_{t}\}$, then in equation (4), the coefficient of $\alpha_{x}$ is exactly 0. If $x\notin\{i,i+1,\ldots,j\}\cup\{\ell_{1},\ell_{2},\ldots,\ell_{t}\}$, then in equation (4), the coefficient of $\alpha_{x}$ is exactly 1. In either case, $\tilde{\alpha}-(\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})-\sum_{h=1}^{t}\alpha_{\ell_{h}}$, can be written as a sum of positive roots. Therefore, $s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}\in\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$. ∎
Having characterized the elements of $\mathcal{A}(\tilde{\alpha},\alpha_{i}+\cdots+\alpha_{j})$ for all $1\leq i\leq j\leq r$, we now consider the cardinality of the set. To do so requires the following definition and enumerative result.
We begin by defining the sequence of numbers commonly referred to as Fibonacci [13] but was first introduced in the literature by Pingala [5].
Definition 2.2.
For any $n\geq 3$, define the Fibonacci number, denoted $F_{n}$, by the recursive formula:
$F_{n}=F_{n-1}+F_{n-2},$
with seed values $F_{1}=F_{2}=1$.
We will utilize the following enumerative result in later results.
Proposition 2.3.
[…]
Proof.
We proceed by induction. For $n=1$, the only options are to select no integers, hence the $\emptyset$, or to select $1$. There are two such choices and note $F_{1+2}=F_{3}=2$, and more over $F_{3}=F_{2}+F_{1}$, as claimed. Assume that for all $k\leq n$ the number of ways to select a subset of $[k]$ consisting of only nonconsecutive integers is given by $F_{k+2}$. Now let $k=n+1$. Note that for any subset of $[n-1]$ consisting of nonconsecutive integers, we can include the integer $n+1$ to get a subset of $[n+1]$ consisting of nonconsecutive integers. Then by induction, there are $F_{n-1+2}=F_{n+1}$ many such subsets, all of which contain the integer $n+1$. Moreover, any subset of $[n]$ consisting of nonconsecutive integers is also a subset of $[n+1]$ with the same property and do not contain $n+1$. By induction hypothesis, there are $F_{n+2}$ such subsets. As these subsets are disjoint (since they either contain $n+1$ or not), then we have shown that the number of subsets of $[n+1]$ consisting of nonconsecutive integers is given by $F_{n+2}+F_{n+1}$, which is equal to $F_{n+3}$, as desired. ∎
Harris in [7] gives another enumerative result which we use in the following section counts subsets of $k$ nonconsecutive integers from the set $[n]$.
Lemma 2.4.
The number of ways to select $k$ nonconsecutive integers from the set $[n]$ is $\binom{n+1-k}{k}.$
Together, Proposition 2.3 and Lemma 2.4 imply the combinatorial identity
$\sum_{k\geq 0}\binom{n+1-k}{k}=F_{n+2}.$
We are now able to enumerate the elements of $\mathcal{A}(\tilde{\alpha},\mu)$ for any $\mu\in\Phi^{+}$.
Corollary 2.5.
If $1\leq i\leq j\leq r$, then
$|\mathcal{A}(\tilde{\alpha},\alpha_{i}+\cdots+\alpha_{j})|=F_{i}\cdot F_{r-j+1},$
[…]
Proof.
Let $1\leq i\leq j\leq r$. By Theorem 2.1, we know $\sigma\in\mathcal{A}(\tilde{\alpha},\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j})$ if and only if $\sigma=1$ or $\sigma=s_{\ell_{1}}\cdots s_{\ell_{k}}s_{\ell_{k+1}}\cdots s_{\ell_{t}}$ for some collection of nonconsecutive integers $12\leq\ell_{1},\ell_{2},\ldots,\ell_{k}\leq i-$ 1 and $j+1\leq\ell_{k+1},\ldots,\ell_{t}\leq r-1$.
Selecting a subset consisting of nonconsecutive integers from the set $\{2,\ldots,i-1\}$ is the same as selecting a subset consisting of nonconsecutve integers from the set $\{1,\ldots,i-2\}$. By Proposition 2.3, there are $F_{i}$ such subsets. The same process can be done for selecting a subset consisting of nonconsecutve integers from the set $\{j+1,\ldots,r-1\}$ to get $F_{r-j+1}$ such subsets.
The choices for subsets $\{2,\ldots,i-1\}$ and $\{j+1,\ldots,r-1\}$ are disjoint and independent of each other. Thus the total number of ways to select nonconsecutive integers is given by the product $F_{i}\cdot F_{r-j+1}$. Therefore, $|\mathcal{A}(\tilde{\alpha},\alpha_{i}+\cdots+\alpha_{j})|=F_{i}\cdot F_{r-j+1}$. ∎
3 The $q$-Analog of Kostant’s Partition Function
Now we are going to give formulas for the value of the $q$-analog of Kostant’s partition function when the input is $\sigma(\tilde{\alpha}+\rho)-\rho-\mu$ with $\mu\in\Phi^{+}$ and $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$. To begin, we give some counts for the number of $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$ with a fixed length.
Lemma 3.1.
If $\mu=\alpha_{1}+\cdots+\alpha_{j}$ with $11\leq j\leq r-$ 1, then
[…]
|$|\{\sigma\in\mathcal{A}(\lambda,\mu)\;|\;\ell(\sigma)=k\mbox{ and$\sigma$ does not contain $s_{j+1}$}\}|$ $=\binom{r-j-1-k}{k}.$ (6)
If $\mu=\alpha_{i}+\cdots+\alpha_{r}$ with $1<i\leq r$, then
[…]
|$|\{\sigma\in\mathcal{A}(\lambda,\mu)\;|\;\ell(\sigma)=k\mbox{ and$\sigma$ does not contain $s_{i-1}$}\}|$ $=\binom{i-2-k}{k}.$ (8)
Proof.
Let $\mu=\alpha_{1}+\cdots+\alpha_{j}$ with $1\leq j<r$. Let $\sigma$ contain $s_{j+1}$. Then we select $k$ nonconsecutive integers from the set $\{j+3,\cdots,r-1\}$. Re-index the elements to be the set $\{1,\ldots,r-j-3\}$. By Lemma 2.4, we get $\binom{r-j-2-k}{k}$. Thus, |$|\{\sigma\in\mathcal{A}(\lambda,\mu)\;|\;\ell(\sigma)=k\mbox{ and $\sigma$ contains $s_{j+1}$}\}|=\binom{r-j-2-k}{k}$.
Let $\sigma$ not contain the element $s_{j+1}$. Then we select $k$ nonconsecutive integers from the set
$\{j+2,\ldots,r-1\}$. Re-index the elements to get the set $\{1,\ldots,r-j-2\}$. By Lemma 2.4, we get $\binom{r-j-1-k}{k}$. Thus, |$|\{\sigma\in\mathcal{A}(\lambda,\mu)\;|\;\ell(\sigma)=k\mbox{ and $\sigma$ does not contain $s_{j+1}$}\}|=\binom{r-j-1-k}{k}$.
The proof for the case where $\mu=\alpha_{i}+\cdots+\alpha_{r}$ with $1<i\leq r$ is analogous. ∎
Lemma 3.2.
If $\mu=\alpha_{1}+\cdots+\alpha_{j}$ with $11\leq j\leq r-$ 1, then
$\max\{\sigma\in\mathcal{A}(\lambda,\mu)\;|\;\ell(\sigma)=k\mbox{ and $\sigma$ contains $s_{j+1}$}\}$ $=\left\lfloor\frac{r-j-2}{2}\right\rfloor\mbox{ and}$
$\max\{\sigma\in\mathcal{A}(\lambda,\mu)\;|\;\ell(\sigma)=k\mbox{ and $\sigma$ does not contain $s_{j+1}$}\}$ $=\left\lfloor\frac{r-j-1}{2}\right\rfloor$
If $\mu=\alpha_{i}+\cdots+\alpha_{r}$ with $2\leq i\leq r$, then
$\max\{\sigma\in\mathcal{A}(\lambda,\mu)\;|\;\ell(\sigma)=k\mbox{ and $\sigma$ contains $s_{i-1}$}\}$ $=\left\lfloor\frac{i-3}{2}\right\rfloor\mbox{ and}$
$\max\{\sigma\in\mathcal{A}(\lambda,\mu)\;|\;\ell(\sigma)=k\mbox{ and $\sigma$ does not contain $s_{i-1}$}\}$ $=\left\lfloor\frac{i-2}{2}\right\rfloor.$
Proof.
Let $\mu=\alpha_{1}+\cdots+\alpha_{j}$ with $11\leq j\leq r-$ 1 and $\ell(\sigma)=k$. Suppose $\sigma$ contains $s_{j+1}$. Then the smallest integer greater than and nonconsecutive to $j+1$ that can be selected is $j+3$. Then we select $k$ nonconsecutive integers from the set $\{j+3,\ldots,r-1\}$, which is the same count as if we select $k$ nonconsecutive integers from the set to get $\{1,\ldots,r-j-3\}$. The maximum number of nonconsecutive numbers that can be selected from the set $\{1,\ldots,r-j-3\}$ is $\left\lfloor\frac{r-j-2}{2}\right\rfloor$.
Suppose $\sigma$ does not contains $s_{j+1}$. Then the smallest integer greater than and nonconsecutive to $j+1$ that can be selected is $j+2$. Then we select $k$ nonconsecutive integers from the set $\{j+2,\ldots,r-1\}$, which is the same count as if we select $k$ nonconsecutive integers from the set $\{1,\ldots,r-j-2\}$. The maximum number of nonconsecutive numbers that can be selected from the set $\{1,\ldots,r-j-2\}$ is $\left\lfloor\frac{r-j-1}{2}\right\rfloor$.
The case where $\mu=\alpha_{i}+\cdots+\alpha_{r}$ with $2\leq i\leq r$ and $\ell(\sigma)=k$ is analogous.∎
Definition 3.3.
For any weight $\xi=c_{1}\alpha_{1}+c_{2}\alpha_{2}+\cdots+c_{r}\alpha_{r}$ with $c_{i}\in\mathbb{N}$, let $h(\xi)=\sum_{i=1}^{r}c_{i}$ denote the height of $\xi$.
The following technical lemma helps in giving formulas for the value of the $q$-analog.
Lemma 3.4.
Let $s,t\geq 1$. If $\beta=\alpha_{1}+\alpha_{2}+\cdots+\alpha_{s}$ and $\gamma=\alpha_{t}+\alpha_{t+1}+\cdots+\alpha_{s+t-1}$, then
$\wp_{q}(\beta)=\wp_{q}(\gamma)=q(1+q)^{s-1}.$
Proof.
Let $\beta=\alpha_{1}+\alpha_{2}+\cdots+\alpha_{s}$ and $\gamma=\alpha_{t}+\alpha_{t+1}+\cdots+\alpha_{s+t-1}$. Notice that $h(\beta)=s$ and $h(\gamma)=(s+t-1)-t+1=s$. By [6, Proposition 3.1], we know
$\wp_{q}(\beta)=\wp_{q}(\alpha_{1}+\alpha_{2}+\cdots+\alpha_{s})=q(1+q)^{s-1}.$
Now we compute $\wp_{q}(\gamma)$. In general, we can write $\gamma$ as a sum of $y\in[s]$ many positive roots by selecting where to place the parentheses enclosing consecutively indexed positive roots. This can be done in $\binom{s-1}{y-1}$ ways. Thus, we get
$\wp_{q}(\gamma)=q^{s}+(s-1)q^{s-1}+\binom{s-1}{2}q^{s-2}+\cdots+\binom{s-1}{y-1}q^{y}+\cdots+(s-1)q^{2}+q$ $=\sum_{y=1}^{s}\binom{s-1}{y-1}q^{y}.$
Reindex $y$ to be $y+1$ and use the Binomial Theorem to get
$\sum_{y=0}^{s-1}\binom{s-1}{y}q^{y+1}=q\sum_{y=0}^{s-1}\binom{s-1}{y}q^{y}=q(1+q)^{s-1}.$
Therefore, $\wp_{q}(\beta)=\wp_{q}(\gamma)$. ∎
Notice that if the height of $\beta$ and $\gamma$ are the same, then by Lemma 3.4, the resulting polynomials $\wp_{q}(\beta)$ and $\wp_{q}(\gamma)$ are the same. Namely, reindexing a positive root of the same height does not change the value of the $q$-analog on it.
The following propositions give the value of the $q$-analog of Kostant’s partition function when the input is $\sigma(\tilde{\alpha}+\rho)-\rho-\mu$, with $\mu\in\Phi^{+}$ and $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$.
Proposition 3.5.
Let $\mu=\alpha_{1}+\cdots+\alpha_{r}$. If $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$, then $\wp_{q}(\sigma(\lambda+\rho)-\rho-\mu)=1$.
Proof.
By Theorem 2.1, in order for $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$, it must be that $\sigma=s_{\ell_{1}}s_{\ell_{2}}\cdots s_{\ell_{a}}s_{\ell_{a+1}}s_{\ell_{a+2}}\cdots s_{\ell_{k}}$ where $12\leq\ell_{1}<\ell_{2}<\cdots<\ell_{a}\leq i-$ 1 and $j+1\leq\ell_{a+1}<\ell_{a+2}<\cdots<\ell_{k}\leq r-1$ are all nonconsecutive integers. We cannot use $s_{1},\ldots,s_{r}$ in any subword of $\sigma$. This implies that $\sigma=1$ and $\mathcal{A}(\tilde{\alpha},\mu)=\{1\}$. Hence $\wp_{q}(\sigma(\lambda+\rho)-\rho-\mu)=\wp_{q}(1(\tilde{\alpha}+\rho)-\rho-(\tilde{\alpha}))=\wp_{q}(0)=1$, as desired. ∎
Proposition 3.6.
  1. (1)
    If $\mu=\alpha_{1}+\cdots+\alpha_{j}$, where $11\leq j\leq r-$ 1 and if $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$, then
    $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ =$=\begin{cases}q^{\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)}&\mbox{if $s_{j+1}$ is contained in $\sigma$}\\ q^{\ell(\sigma)+1}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}&\mbox{otherwise.}\end{cases}$ (9)
  2. (2)
    If $\mu=\alpha_{i}+\cdots+\alpha_{r}$, where $2\leq i\leq r$, and if $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$, then
    $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ =$=\begin{cases}q^{\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)}&\mbox{if $s_{i-1}$ is containted in $\sigma$}\\ q^{\ell(\sigma)+1}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}&\mbox{otherwise.}\end{cases}$ (10)
Proof.
Let $\mu=\alpha_{1}+\cdots+\alpha_{j}$. By Theorem 2.1, if $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$, then $\sigma=s_{\ell_{1}}s_{\ell_{2}}\cdots s_{\ell_{k}}$ where
$j+1\leq\ell_{1}<\ell_{2}<\cdots<\ell_{k}\leq r-1$ are all nonconsecutive integers.
Suppose $\ell_{1}=j+1$, i.e. $s_{j+1}$ is contained in $\sigma$. Then take $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-(\alpha_{1}+\cdots+\alpha_{j}))$ and simplify $\sigma(\tilde{\alpha}+\rho)-\rho-(\alpha_{1}+\cdots+\alpha_{j})$ by segmenting the sum of $\alpha$’s by indicating where there was a $s_{x}$ in $\sigma$. This yields
$\wp_{q}([\alpha_{\ell_{1}+1}+\cdots+\alpha_{\ell_{2}-1}]+[\alpha_{\ell_{2}+1}+\cdots+\alpha_{\ell_{3}-1}]+\cdots+[\alpha_{\ell_{k}+1}+\cdots+\alpha_{r}]).$ (11)
Substituting $\ell_{1}=j+1$ into equation (11) yields
$\wp_{q}([\alpha_{j+2}+\cdots+\alpha_{\ell_{2}-1}]+[\alpha_{\ell_{2}+1}+\cdots+\alpha_{\ell_{3}-1}]+\cdots+[\alpha_{\ell_{k}+1}+\cdots+\alpha_{r}]).$ (12)
Since the subsets of the positive roots used to write each of the sums within the brackets are pairwise disjoint, we can rewrite equation (12) as the product
$\wp_{q}(\alpha_{j+2}+\cdots+\alpha_{\ell_{2}-1})\cdot\wp_{q}(\alpha_{\ell_{2}+1}+\cdots+\alpha_{\ell_{3}-1})\cdots\wp_{q}(\alpha_{\ell_{k}+1}+\cdots+\alpha_{r}).$ (13)
By Lemma 3.4, we get that equation (13) simplifies to
$q(1+q)^{(\ell_{2}-j-2)-1}q(1+q)^{(\ell_{3}-\ell_{2}-1)-1}\cdots q(1+q)^{(r-\ell_{k})-1}$ $=q^{k}(1+q)^{r-j-2(k)}=q^{\ell(\sigma)}(1+q)^{r-j-2\ell(\sigma)}$ (14)
where equation (14) holds by simplifying the exponents for $q$ and $(1+q)$ and noting that there are $k=\ell(\sigma)$ terms.
Now suppose $\ell_{1}\neq j+1$, i.e. $s_{j+1}$ is not contained in $\sigma$. Then take $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-(\alpha_{1}+\cdots+\alpha_{j}))$ and simplify $\sigma(\tilde{\alpha}+\rho)-\rho-(\alpha_{1}+\cdots+\alpha_{j})$ by segmenting the sum of $\alpha$’s by indicating where there was a $s_{x}$ in $\sigma$. This yields
$\wp_{q}([\alpha_{j+1}+\cdots+\alpha_{\ell_{1}-1}]+[\alpha_{\ell_{1}+1}+\cdots+\alpha_{\ell_{2}-1}]+\cdots+[\alpha_{\ell_{k}+1}+\cdots+\alpha_{r}]).$ (15)
Since the subsets of the positive roots used to write each of the sums within the brackets are pairwise disjoint, we can rewrite the equation (15) as the product
$\wp_{q}(\alpha_{j+1}+\cdots+\alpha_{\ell_{1}-1})\cdot\wp_{q}(\alpha_{\ell_{1}+1}+\cdots+\alpha_{\ell_{2}-1})\cdots\wp_{q}(\alpha_{\ell_{k}+1}+\cdots+\alpha_{r}).$ (16)
By Lemma 3.4, we get that equation (16) simplifies to
$q(1+q)^{(\ell_{1}-j-1)-1}\cdot q(1+q)^{(\ell_{2}-\ell_{1}-1)-1}\cdots q(1+q)^{(r-\ell_{k})-1}$ $=q^{k+1}(1+q)^{r-j-2(k)-1}$
$=q^{\ell(\sigma)+1}(1+q)^{r-j-2\ell(\sigma)-1}$ (17)
where equation (17) holds by simplifying the exponents for $q$ and $(1+q)$.
Thus, since $h(\mu)=j$, we arrive at
$\wp_{q}(\sigma(\lambda+\rho)-\rho-\mu)=\begin{cases}q^{\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)}&\mbox{if $\ell_{1}=j+1$}\\ q^{\ell(\sigma)+1}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}&\mbox{if $\ell_{1}\neq j+1$}.\end{cases}$
The proof for statement (2) is analogous. ∎
Proposition 3.7.
Let $\mu=\alpha_{i}+\cdots+\alpha_{j}$, where $12\leq i\leq j\leq r-$ 1. If $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$, then
$\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ =$=\begin{cases}q^{2+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-2}&\mbox{if $s_{i-1}$ and $s_{j+1}$ are not containted in $\sigma$}\\ q^{1+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}&\mbox{if $s_{i-1}$ is contained in $\sigma$ but $s_{j+1}$ is not}\\ q^{1+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}&\mbox{if $s_{i-1}$ is not contained in $\sigma$ but $s_{j+1}$ is}\\ q^{\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)}&\mbox{if $s_{i-1}$ and $s_{j+1}$ are contained in $\sigma$.}\end{cases}$ (18)
Proof.
Let $\mu=\alpha_{i}+\cdots+\alpha_{j}$, where $12\leq i\leq j\leq r-$ 1 and note $h(\mu)=j-i+1$.
By Theorem 2.1, since $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$, then $\sigma=s_{\ell_{1}}s_{\ell_{2}}\cdots s_{\ell_{a}}s_{\ell_{a+1}}s_{\ell_{a+2}}\cdots s_{\ell_{k}}$ where
$12\leq\ell_{1}<\ell_{2}<\cdots<\ell_{a}\leq i-$ 1 and $j+1\leq\ell_{a+1}<\ell_{a+2}<\cdots<\ell_{k}\leq r-1$ are all nonconsecutive integers. Let $\sigma=\sigma_{1}\sigma_{2}$ where $\sigma_{1}=s_{\ell_{1}}s_{\ell_{2}}\cdots s_{\ell_{a}}$ and $\sigma_{2}=s_{\ell_{a+1}}s_{\ell_{a+2}}\cdots s_{\ell_{k}}$. Since the $\ell_{x}$’s appearing in $\sigma$ are nonconsecutive, then the $s_{\ell_{x}}$’s can appear in $\sigma$ in any order. This implies $\sigma_{1}\sigma_{2}=\sigma_{2}\sigma_{1}$. Then
$\wp_{q}(\sigma(\lambda+\rho)-\rho-\mu)$ $=\wp_{q}\left(\tilde{\alpha}+\rho-\left(\sum_{s_{\ell_{b}}\in\sigma_{1}}\alpha_{\ell_{b}}\right)-\left(\sum_{s_{\ell_{d}}\in\sigma_{2}}\alpha_{\ell_{d}}\right)-\rho-\mu\right)$
$=\wp_{q}\left(\sum_{k=1}^{r}\alpha_{k}-\left(\sum_{s_{\ell_{b}}\in\sigma_{1}}\alpha_{\ell_{b}}\right)-\left(\sum_{s_{\ell_{d}}\in\sigma_{2}}\alpha_{\ell_{d}}\right)-\sum_{k=i}^{j}\alpha_{k}\right).$
Simplifying $\sum_{k=1}^{r}\alpha_{k}-\left(\sum_{s_{\ell_{b}}\in\sigma_{1}}\alpha_{\ell_{b}}\right)-\left(\sum_{s_{\ell_{d}}\in\sigma_{2}}\alpha_{\ell_{d}}\right)-\sum_{k=i}^{j}\alpha_{k}$ by segmenting the sums and indicating where there was a $s_{x}$ in $\sigma$ yields
$\wp_{q}([\alpha_{1}+\cdots+\alpha_{\ell_{1}-1}]+\cdots+[\alpha_{\ell_{a}+1}+\cdots+\alpha_{i-1}]+[\alpha_{j+1}+\cdots+\alpha_{\ell_{a+1}-1}]+\cdots+[\alpha_{\ell_{k}+1}+\cdots+\alpha_{r}]).$
Since the subsets of the positive roots used to write each of the sums within the brackets are pairwise disjoint, we can rewrite the equation as the product
$\wp_{q}$ $([\alpha_{1}+\cdots+\alpha_{\ell_{1}-1}]+\cdots+[\alpha_{\ell_{a}+1}+\cdots+\alpha_{i-1}])\cdot\wp_{q}([\alpha_{j+1}+\cdots+\alpha_{\ell_{a+1}-1}]+\cdots+[\alpha_{\ell_{k}+1}+\cdots+\alpha_{r}])$
$=\wp_{q}(\sigma_{1}(\tilde{\alpha}+\rho)-\rho-(\alpha_{i}+\cdots+\alpha_{r}))\cdot\wp_{q}(\sigma_{2}(\tilde{\alpha}+\rho)-\rho-(\alpha_{1}+\cdots+\alpha_{j})).$ (19)
Suppose $s_{i-1}$ and $s_{j+1}$ are not contained in $\sigma$. Proposition 3.6 allows us to simplify equation (19) to get
$\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ $=q^{\ell(\sigma_{1})+1}(1+q)^{i-2-2\ell(\sigma_{1})}q^{\ell(\sigma_{2})+1}(1+q)^{r-j-2\ell(\sigma)-1}$
$=q^{2+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-2}.$ (20)
Equation (20) is achieved by combining the exponents for the $q$’s and the $(1+q)$’s and by recalling the fact that $\ell(\sigma_{1})+\ell(\sigma_{2})=\ell(\sigma)$ and $h(\mu)=j-i+1$.
Suppose $s_{i-1}$ is contained in $\sigma$ but $s_{j+1}$ is not contained in $\sigma$. Proposition 3.6 allows us to simplify equation (19) to get
$\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ $=q^{\ell(\sigma_{1})}(1+q)^{i-1-2\ell(\sigma_{1})}q^{\ell(\sigma_{2})+1}(1+q)^{r-j-2\ell(\sigma)-1}$
$=q^{1+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}.$ (21)
Equation (21) is achieved by combine the exponents for the $q$’s and the $(1+q)$’s and by recalling the fact that $\ell(\sigma_{1})+\ell(\sigma_{2})=\ell(\sigma)$ and $h(\mu)=j-i+1$.
The case where $s_{i-1}$ is not contained in $\sigma$ and $s_{j+1}$ is contained in $\sigma$, is analogous to the case where $s_{i-1}$ is contained in $\sigma$ but $s_{j+1}$ is not contained in $\sigma$, and hence we omit the argument.
Suppose $s_{i-1}$ and $s_{j+1}$ are contained in $\sigma$. Proposition 3.6 allows us to simplify equation (19) to get
$\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ $=q^{\ell(\sigma_{1})}(1+q)^{(i-1)-2\ell(\sigma_{1})}q^{\ell(\sigma_{2})}(1+q)^{(r-j)-2\ell(\sigma_{2})}$
$=q^{\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)}.$ (22)
Equation (22) is achieved by combining the exponents for the $q$’s and the $(1+q)$’s and by recalling the fact that $\ell(\sigma_{1})+\ell(\sigma_{2})=\ell(\sigma)$ and $h(\mu)=j-i+1$.
Therefore,
$\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)=\begin{cases}q^{2+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-2}&\mbox{if $s_{i-1}$ and $s_{j+1}$ are not containted in $\sigma$}\\ q^{1+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}&\mbox{if $s_{i-1}$ is contained in $\sigma$ but $s_{j+1}$ is not}\\ q^{1+\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)-1}&\mbox{if $s_{i-1}$ is not contained in $\sigma$ but $s_{j+1}$ is}\\ q^{\ell(\sigma)}(1+q)^{r-h(\mu)-2\ell(\sigma)}&\mbox{if $s_{i-1}$ and $s{j+1}$are contained in $\sigma$.}\end{cases}$
4 The $q$-Multiplicity of a Positive Root is a Power of $q$
We have all the information needed, we are now ready to prove Theorem 4.1, which we restate for ease of reference.
Theorem 4.1.
Fix $1\leq i\leq j\leq r$ and let $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}\in\Phi^{+}$. Then
$m_{q}(\tilde{\alpha},\mu)=q^{r-h(\mu)},$
where $h(\mu)=j-i+1$ denotes the height of $\mu$.
Proof.
Let $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}\in\Phi^{+}$ with $1\leq i\leq j\leq r$, which has height $h(\mu)=j-i+1$. By definition
$m_{q}(\tilde{\alpha},\mu)=\sum_{\sigma\in W}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu).$
By Theorem 2.1, we can reduce the computing to be only on the Weyl alternation set. Namely,
$m_{q}(\tilde{\alpha},\mu)=\sum_{\sigma\in\mathcal{A}(\tilde{\alpha},\mu)}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu).$
For any $x\in[r]$, let
  • $N_{x}=\{\sigma\in\mathcal{A}(\tilde{\alpha},\mu)\;|\;\mbox{$\sigma$ does not contain $s_{x}$ in its reduced word}\}$
  • $Y_{x}=\{\sigma\in\mathcal{A}(\tilde{\alpha},\mu)\;|\;\mbox{$\sigma$ contains $s_{x}$ in its reduced word}\}$.
For each case, we partition the set $\mathcal{A}(\tilde{\alpha},\mu)$ into disjoint subsets of $N_{x}$ and $Y_{x}$ according to the construction of $\mu$.
Case 1: Fix $\mu=\alpha_{i}+\cdots+\alpha_{j}$ where $12\leq i\leq j\leq r-$ 1. To begin, we follow the same construction of equation (19), which we restate here:
$\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)=\wp_{q}(\sigma_{1}(\tilde{\alpha}+\rho)-\rho-(\alpha_{i}+\cdots+\alpha_{r}))\cdot\wp_{q}(\sigma_{2}(\tilde{\alpha}+\rho)-\rho-(\alpha_{1}+\cdots+\alpha_{j}))$ (23)
where $\ell(\sigma)=k$, $\sigma=\sigma_{1}\sigma_{2}$, $\sigma_{1}=s_{\ell_{1}}s_{\ell_{2}}\cdots s_{\ell_{a}}$ where $12\leq\ell_{1}<\ell_{2}<\cdots<\ell_{a}\leq i-$ 1, and $\sigma_{2}=s_{\ell^{\prime}_{1}}s_{\ell^{\prime}_{2}}\cdots s_{\ell^{\prime}_{b}}$ where $j+1\leq\ell^{\prime}_{1}<\ell^{\prime}_{2}<\cdots<\ell^{\prime}_{b}\leq r-1$ are all nonconsecutive integers. Note $\ell(\sigma_{1})=a$, $\ell(\sigma_{2})=b$, and $\ell(\sigma)=k=a+b$. By construction, we have
$\mathcal{A}(\tilde{\alpha},\mu)=(N_{i-1}\cap N_{j+1})\sqcup(N_{i-1}\cap Y_{j+1})\sqcup(Y_{i-1}\cap N_{j+1})\sqcup(Y_{i-1}\cap Y_{j+1})$ (24)
where $\sigma$ is considered to be in the set $N_{i-1}\cap N_{j+1}$, $N_{i-1}\cap Y_{j+1}$, $Y_{i-1}\cap N_{j+1}$, or $Y_{i-1}\cap Y_{j+1}$.
  • Let $\sigma\in N_{i-1}\cap N_{j+1}$ such that $\ell(\sigma)=\ell(\sigma_{1})+\ell(\sigma_{2})=k$ where we let $\ell(\sigma_{1})=a$ and $\ell(\sigma_{2})=b$. Then
    $\wp_{q}(\sigma_{1}(\tilde{\alpha}+\rho)-\rho-(\alpha_{i}+\cdots+\alpha_{r}))$ $=\wp_{q}\left(\tilde{\alpha}+\rho-\sum_{s_{x}\in\sigma_{1}}\alpha_{x}-\rho-(\alpha_{i}+\cdots+\alpha_{r})\right)$
    $=\wp_{q}\left(\alpha_{1}+\cdots+\alpha_{i-1}+\gamma-\sum_{s_{x}\in\sigma_{1}}\alpha_{x}-\gamma\right)$
    $=\wp_{q}(\sigma_{1}(\beta+\gamma)-\gamma)$
    where $\beta$ is the highest root of $\mathfrak{sl}_{i}(\mathbb{C})$, $\gamma$ is the sum of the fundamental weights of $\mathfrak{sl}_{i}(\mathbb{C})$, and $\sigma_{1}\in\mathcal{A}(\beta,0)$ [6, Theorem 2.1].
    Similarly,
    $\wp_{q}(\sigma_{2}(\tilde{\alpha}+\rho)-\rho-(\alpha_{1}+\cdots+\alpha_{j}))=\wp_{q}(\sigma_{2}(\beta^{\prime}+\gamma^{\prime})-\gamma^{\prime})$
    where $\beta^{\prime}$ is the highest root of $\mathfrak{sl}_{r-j+1}(\mathbb{C})$, $\gamma^{\prime}$ is the sum of the fundamental weights of $\mathfrak{sl}_{r-j+1}(\mathbb{C})$, and $\sigma_{2}\in\mathcal{A}(\beta^{\prime},0)$ [6, Theorem 2.1].
    By [6, Proposition 3.2], note that
    $\wp_{q}(\sigma_{1}(\beta+\gamma)-\gamma)$ $=q^{1+\ell(\sigma_{1})}(1+q)^{(i-1)-1-2\ell(\sigma_{1})}\mbox{ and}$
    $\wp_{q}(\sigma_{2}(\beta^{\prime}+\gamma^{\prime})-\gamma^{\prime})$ $=q^{1+\ell(\sigma_{2})}(1+q)^{(r-j)-1-2\ell(\sigma_{2})}.$
    Given the equality in equation (23), note that
    $\sum_{\sigma\in N_{i-1}\cap N_{j+1}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$
    $~{}=\left(\sum_{\sigma_{1}\in\mathcal{A}(\beta,0)}(-1)^{\ell(\sigma_{1})}\wp_{q}(\sigma_{1}(\beta+\gamma)-\gamma)\right)\cdot\left(\sum_{\sigma_{2}\in\mathcal{A}(\beta^{\prime},0)}(-1)^{\ell(\sigma_{2})}\wp_{q}(\sigma_{2}(\beta^{\prime}+\gamma^{\prime})-\gamma^{\prime})\right)$
    $=m_{q}(\beta,0)\cdot m_{q}(\beta^{\prime},0)$
    $=\left(\sum_{t=1}^{i-1}q^{t}\right)\left(\sum_{t=1}^{r-j}q^{t}\right),$ (25)
    where the equality in (25) follows from [2, Theorem 6.1].
  • If $\sigma\in N_{i-1}\cap Y_{j+1}$ such that $\ell(\sigma)=\ell(\sigma_{1})+\ell(\sigma_{2})=k+1$ where we let $\ell(\sigma_{1})=a$ and $\ell(\sigma_{2})=b+1$. Then $\wp_{q}(\sigma_{1}(\tilde{\alpha}+\rho)-\rho-(\alpha_{i}+\cdots+\alpha_{r}))=\wp_{q}(\sigma_{1}(\beta+\gamma)-\gamma)$ like above. However,
    $\wp_{q}(\sigma_{2}(\tilde{\alpha}+\rho)-\rho-(\alpha_{1}+\cdots+\alpha_{j}))=\wp_{q}(\sigma_{2}(\beta^{\prime}+\gamma^{\prime})-\gamma^{\prime})$
    where $\beta^{\prime}$ is the highest root of $\mathfrak{sl}_{r-j}(\mathbb{C})$, $\gamma^{\prime}$ is the sum of the fundamental weights of $\mathfrak{sl}_{r-j}(\mathbb{C})$, and $\sigma_{2}\in\mathcal{A}(\beta^{\prime},0)$ [6, Theorem 2.1].
    By [6, Proposition 3.2], note that
    $\wp_{q}(\sigma_{1}(\beta+\gamma)-\gamma)$ $=q^{1+\ell(\sigma_{1})}(1+q)^{(i-1)-1-2\ell(\sigma_{1})}\mbox{ and }$
    $\wp_{q}(\sigma_{2}(\beta^{\prime}+\gamma^{\prime})-\gamma^{\prime})$ $=q^{1+\ell(\sigma_{2})}(1+q)^{(r-j-1)-1-2\ell(\sigma_{2})}.$
    Given the equality in equation (23), note that
    $\sum_{\sigma\in N_{i-1}\cap Y_{j+1}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$
    $~{}=\left(\sum_{\sigma_{1}\in\mathcal{A}(\beta,0)}(-1)^{\ell(\sigma_{1})}\wp_{q}(\sigma_{1}(\beta+\gamma)-\gamma)\right)\cdot\left(\sum_{\sigma_{2}\in\mathcal{A}(\beta^{\prime},0)}(-1)^{\ell(\sigma_{2})}\wp_{q}(\sigma_{2}(\beta^{\prime}+\gamma^{\prime})-\gamma^{\prime})\right)$
    $=m_{q}(\beta,0)\cdot(-1)m_{q}(\beta^{\prime},0)$ (26)
    where the equality in (26) is satisfied because $\ell(\sigma_{2})=b+1$. Then we get
    $(-1)\left(\sum_{t=1}^{i-1}q^{t}\right)\left(\sum_{t=1}^{r-j-1}q^{t}\right),$ (27)
    where expression (27) follows from [2, Theorem 6.1].
  • If $\sigma\in Y_{i-1}\cap N_{j+1}$ such that $\ell(\sigma)=\ell(\sigma_{1})+\ell(\sigma_{2})=k+1$ where we let $\ell(\sigma_{1})=a+1$ and $\ell(\sigma_{2})=b$. Now
    $\wp_{q}(\sigma_{1}(\tilde{\alpha}+\rho)-\rho-(\alpha_{i}+\cdots+\alpha_{r}))=\wp_{q}(\sigma_{1}(\beta+\gamma)-\gamma)$
    where $\beta$ is the highest root of $\mathfrak{sl}_{i-1}(\mathbb{C})$, $\gamma$ is the sum of the fundamental weights of $\mathfrak{sl}_{i-1}(\mathbb{C})$, and $\sigma_{1}\in\mathcal{A}(\beta,0)$ [6, Theorem 2.1].
    Like above, $\wp_{q}(\sigma_{2}(\tilde{\alpha}+\rho)-\rho-(\alpha_{1}+\cdots+\alpha_{j}))=\wp_{q}(\sigma_{2}(\beta^{\prime}+\gamma^{\prime})-\gamma^{\prime})$.
    By [6, Proposition 3.2] note that
    $\wp_{q}(\sigma_{1}(\beta+\gamma)-\gamma)$ $=q^{1+\ell(\sigma_{1})}(1+q)^{(i-2)-1-2\ell(\sigma_{1})}\mbox{ and}$
    $\wp_{q}(\sigma_{2}(\beta^{\prime}+\gamma^{\prime})-\gamma^{\prime})$ $=q^{1+\ell(\sigma_{2})}(1+q)^{(r-j)-1-2\ell(\sigma_{2})}.$
    Given the equality in equation (23), note that
    $\sum_{\sigma\in Y_{i-1}\cap N_{j+1}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$
    $=\left(\sum_{\sigma_{1}\in\mathcal{A}(\beta,0)}(-1)^{\ell(\sigma_{1})}\wp_{q}(\sigma_{1}(\beta+\gamma)-\gamma)\right)\cdot\left(\sum_{\sigma_{2}\in\mathcal{A}(\beta^{\prime},0)}(-1)^{\ell(\sigma_{2})}\wp_{q}(\sigma_{2}(\beta^{\prime}+\gamma^{\prime})-\gamma^{\prime})\right)$
    $=(-1)m_{q}(\beta,0)\cdot m_{q}(\beta^{\prime},0)$ (28)
    where the equality in (28) is satisfied because $\ell(\sigma_{1})=a+1$. Then we get
    $(-1)\left(\sum_{t=1}^{i-2}q^{t}\right)\left(\sum_{t=1}^{r-j}q^{t}\right),$ (29)
    where expression (29) follows from [2, Theorem 6.1].
  • If $\sigma\in Y_{i-1}\cap Y_{j+1}$ such that $\ell(\sigma)=\ell(\sigma_{1})+\ell(\sigma_{2})=k+2$ where we let $\ell(\sigma_{1})=a+1$ and $\ell(\sigma_{2})=b+1$. Like in the above respective corresponding cases,
    $\wp_{q}(\sigma_{1}(\tilde{\alpha}+\rho)-\rho-(\alpha_{i}+\cdots+\alpha_{r}))=\wp_{q}(\sigma_{1}(\beta+\gamma)-\gamma)$ and
    $\wp_{q}(\sigma_{2}(\tilde{\alpha}+\rho)-\rho-(\alpha_{1}+\cdots+\alpha_{j}))=\wp_{q}(\sigma_{2}(\beta^{\prime}+\gamma^{\prime})-\gamma^{\prime})$.
    By [6, Proposition 3.2] note that
    $\wp_{q}(\sigma_{1}(\beta+\gamma)-\gamma)$ $=q^{1+\ell(\sigma_{1})}(1+q)^{(i-2)-1-2\ell(\sigma_{1})},\mbox{ and}$
    $\wp_{q}(\sigma_{2}(\beta^{\prime}+\gamma^{\prime})-\gamma^{\prime})$ $=q^{1+\ell(\sigma_{2})}(1+q)^{(r-j-1)-1-2\ell(\sigma_{2})}.$
    Given the equality in equation (23), note that
    $\sum_{\sigma\in Y_{i-1}\cap Y_{j+1}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$
    $=\left(\sum_{\sigma_{1}\in\mathcal{A}(\beta,0)}(-1)^{\ell(\sigma_{1})}\wp_{q}(\sigma_{1}(\beta+\gamma)-\gamma)\right)\cdot\left(\sum_{\sigma_{2}\in\mathcal{A}(\beta^{\prime},0)}(-1)^{\ell(\sigma_{2})}\wp_{q}(\sigma_{2}(\beta^{\prime}+\gamma^{\prime})-\gamma^{\prime})\right)$
    $=(-1)m_{q}(\beta,0)\cdot(-1)m_{q}(\beta^{\prime},0)$ (30)
    where the equality in (30) is satisfied because $\ell(\sigma_{1})=a+1$ and $\ell(\sigma_{1})=b+1$. Then we get
    $\left(\sum_{t=1}^{i-2}q^{t}\right)\left(\sum_{t=1}^{r-j-1}q^{t}\right),$ (31)
    where expression (31) follows from [2, Theorem 6.1].
Now that we have all cases, we can construct the $m_{q}(\tilde{\alpha},\mu)$ from (24) by adding expression (25), expression (27), expression (29), and expression (31) to get the following:
$m_{q}(\tilde{\alpha},\mu)$ $=\left(\sum_{t=1}^{i-1}q^{t}\right)\left(\sum_{t=1}^{r-j}q^{t}\right)-\left(\sum_{t=1}^{i-1}q^{t}\right)\left(\sum_{t=1}^{r-j-1}q^{t}\right)-\left(\sum_{t=1}^{i-2}q^{t}\right)\left(\sum_{t=1}^{r-j}q^{t}\right)+\left(\sum_{t=1}^{i-2}q^{t}\right)\left(\sum_{t=1}^{r-j-1}q^{t}\right)$
$=q^{r-(j-i+1)}.$
Thus, $m_{q}(\tilde{\alpha},\mu)=q^{r-h(\mu)}$ when $\mu=\alpha_{i}+\cdots+\alpha_{j}\in\Phi^{+}$ with $12\leq i\leq j\leq r-$ 1 and $h(\mu)=j-i+1$.
Case 2: Suppose $\mu=\alpha_{1}+\cdots+\alpha_{j}\in\Phi^{+}$ where $11\leq j\leq r-$ 1. Notice that $h(\mu)=j$.
Observe that $\mathcal{A}(\tilde{\alpha},\mu)=N_{j+1}\sqcup Y_{j+1}$ is a disjoint union. Thus,
$m_{q}(\tilde{\alpha},\mu)=$ $\sum_{\sigma\in{N_{j+1}}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)+\sum_{\sigma\in{Y_{j+1}}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu).$ (32)
First, consider $\sum_{\sigma\in{N_{j+1}}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$. By Lemma 3.2, we know the maximum number of $\sigma\in\mathcal{A}(\tilde{\alpha},\mu)$. By Lemma 3.1, we know there are $\binom{r-j-1-k}{k}$ elements $\sigma\in N_{j+1}$ of length $k$ and that the maximum length of any $\sigma\in N_{j+1}$ is $\left\lfloor\frac{r-j-1}{2}\right\rfloor$. Proposition 3.6 gives the value of the $q$-analog $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ for $\sigma\in N_{j+1}$. Thus, if $\sigma\in N_{j+1}$ and we use [6, Proposition 3.2], then the first sum in equation (32) becomes
$\sum_{k=0}^{\left\lfloor\frac{r-j-1}{2}\right\rfloor}(-1)^{k}\binom{r-j-1-k}{k}q^{k+1}(1+q)^{r-j-1-2k}=\sum_{t=1}^{r-j}q^{t}.$ (33)
Similarly, consider $\sum_{\sigma\in{Y_{j+1}}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$. By Lemma 3.2, we know there are $\binom{r-j-2-k}{k}$ elements $\sigma\in Y_{j+1}$ of length $k$ and that the maximum length of any $\sigma\in Y_{j+1}$ is $\left\lfloor\frac{r-j-2}{2}\right\rfloor$. Proposition 3.6 gives the value of the $q$-analog $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ for $\sigma\in Y_{j+1}$. Thus, if $\sigma\in Y_{j+1}$ and we use [6, Proposition 3.2], then the second sum in equation (32) becomes
$\sum_{k=0}^{\left\lfloor\frac{r-j-2}{2}\right\rfloor}(-1)^{k+1}\binom{r-j-2-k}{k}q^{k+1}(1+q)^{r-j-2(k+1)}=-\sum_{t=1}^{r-j-1}q^{t}.$ (34)
Calculate $m_{q}(\tilde{\alpha},\mu)$ by adding expression (33) and expression (34) to get
$\sum_{t=1}^{r-j}q^{t}-\sum_{t=1}^{r-j-1}q^{t}=q^{r-j}.$ (35)
Thus, $m_{q}(\tilde{\alpha},\mu)=q^{r-h(\mu)}$ since $h(\mu)=j$.
Case 3: Suppose $\mu=\alpha_{i}+\cdots+\alpha_{r}\in\Phi^{+}$ where $2\leq i\leq r$. Notice that $h(\mu)=r-i+1$.
Observe that $\mathcal{A}(\tilde{\alpha},\mu)=N_{i-1}\sqcup Y_{i-1}$ is a disjoint union. Thus,
$m_{q}(\tilde{\alpha},\mu)=\sum_{\sigma\in N_{i-1}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)+\sum_{\sigma\in Y_{i-1}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu).$ (36)
First, consider $\sum_{\sigma\in N_{i-1}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$. By Lemma 3.2, we know the maximum length of an element $\sigma\in N_{i-1}$ is $\left\lfloor\frac{i-2}{2}\right\rfloor$. By Lemma 3.1, we know there are $\binom{i-2-k}{k}$ elements $\sigma\in N_{i-1}$. Proposition 3.6 gives the value of the $q$-analog $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ for $\sigma\in N_{i-1}$. Thus if $\sigma\in N_{i-1}$ and we use [6, Proposition 3.2], the first sum in equation (36) becomes
$\sum_{k=0}^{\left\lfloor\frac{i-2}{2}\right\rfloor}(-1)^{k}\binom{i-2-k}{k}q^{k+1}(1+q)^{i-2-2k}=\sum_{t=1}^{i-1}q^{t}.$ (37)
Similarly, consider $\sum_{\sigma\in Y_{i-1}}(-1)^{\ell(\sigma)}\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$. By Lemma 3.2, we know the maximum length of an element $\sigma\in Y_{i-1}$ is $\left\lfloor\frac{i-3}{2}\right\rfloor$. By Lemma 3.1, we know there are $\binom{i-3-k}{k}$ elements $\sigma\in Y_{i-1}$. Proposition 3.6 gives the value of the $q$-analog $\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)$ for $\sigma\in Y_{i-1}$. Thus if $\sigma\in Y_{i-1}$ and we use [6, Proposition 3.2], the second sum in equation (36) becomes
$\sum_{k=0}^{\left\lfloor\frac{i-3}{2}\right\rfloor}(-1)^{k+1}\binom{i-3-k}{k}q^{k+1}(1+q)^{(i-1)-2{(k+1)}}=-\sum_{t=1}^{i-2}q^{t}.$ (38)
Calculate $m_{q}(\tilde{\alpha},\mu)$ by adding expression (37) and expression (38) to get
$\sum_{t=1}^{i-1}q^{t}-\sum_{t=1}^{i-2}q^{t}=q^{i-1}.$ (39)
Thus, $m_{q}(\tilde{\alpha},\mu)=q^{r-h(\mu)}$ since $h(\mu)=r-i+1$.
Case 4: Suppose $\mu=\alpha_{1}+\cdots+\alpha_{r}\in\Phi^{+}$. Notice that $h(\mu)=r-1+1=r$. Observe that the elements of $\mathcal{A}(\tilde{\alpha},\mu)=\{1\}$. Thus,
$m_{q}(\tilde{\alpha},\mu)=\wp_{q}(\sigma(\tilde{\alpha}+\rho)-\rho-\mu)=\wp_{q}(1(\tilde{\alpha}+\rho)-\rho-\tilde{\alpha})=\wp_{q}(0)=1=q^{r-(r-1+1)}=q^{r-h(\mu)}.$
Therefore, when $r\geq 1$ and $\tilde{\alpha}$ is the highest root of $\mathfrak{sl}_{r+1}(\mathbb{C})$ and $\mu=\alpha_{i}+\alpha_{i+1}+\cdots+\alpha_{j}\in\Phi^{+}$ with $h(\mu)=j-i+1$, then $m_{q}(\tilde{\alpha},\mu)=q^{r-h(\mu)}$. ∎